# CoSpaDi: Compressing LLMS via Calibration-Guided Sparse Dictionary Learning

### v2

@@ Abstract
Post-training compression of large language models (LLMs) largely relies on low-rank weight approximation, which represents each column of a weight matrix in a shared low-dimensional subspace. While this is a computationally efficient strategy, the imposed structural constraint is rigid and can lead to a noticeable model accuracy drop. In this work, we propose CoSpaDi (Compression via Sparse Dictionary Learning), a novel training-free compression framework that replaces low-rank decomposition with a more flexible structured sparse factorization in which each weight matrix is represented with a dense dictionary and a column-sparse coefficient matrix. This formulation enables a union-of-subspaces representation: different columns of the original weight matrix are approximated in distinct subspaces spanned by adaptively selected dictionary atoms, offering greater expressiveness than a single invariant basis. Crucially, CoSpaDi leverages a small calibration dataset to optimize the factorization such that the output activations of compressed projection layers closely match those of the original ones, thereby minimizing functional reconstruction error rather than mere weight approximation. This data-aware strategy preserves better model fidelity without any fine-tuning under reasonable compression ratios. Moreover, the resulting structured sparsity allows efficient sparse–dense matrix multiplication and is compatible with post-training quantization for further memory and latency gains. We evaluate CoSpaDi across multiple Llama and Qwen models under per-layer and per-group settings at 205020-50% compression ratios, demonstrating consistent superiority over state-of-the-art data-aware low-rank methods both in accuracy and perplexity. Our results establish structured sparse dictionary learning as a powerful alternative to conventional low-rank approaches for efficient LLM deployment.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of applications, from conversational agents Brown et al. (2020); OpenAI (2023) to general-purpose reasoning systems Touvron et al. (2023a); Anil et al. (2023), owing to their ability to capture long-range dependencies through attention mechanisms Vaswani et al. (2017); Devlin et al. (2019). However, this success entails substantial memory and computational demands during both training and inference, which severely limits their deployment in resource-constrained environments.
Various compression and acceleration techniques have been proposed over recent years to mitigate these challenges, including pruning Frankle & Carbin (2019); Sanh et al. (2020), quantization Dettmers et al. (2022); Xiao et al. (2023); Yao et al. (2022); Dettmers et al. (2023), and weight decomposition/factorization Denton et al. (2014); Hu et al. (2022); Liu et al. (2024); Zhang et al. (2023); Ma et al. (2019); Chen et al. (2018); Hsu et al. (2022); Yu & Wu (2023); Wang et al. (2025a); Han et al. (2015).
Recently, there has been growing interest in effective compression of large models without retraining or fine-tuning, since this is often computationally prohibitive. A prominent line of work leverages low-rank approximations of weight matrices via Singular Value Decomposition (SVD). Early approaches such as DRONE Chen et al. (2021), FWSVD Hsu et al. (2022) and ASVD Yuan et al. (2023) introduced activation-aware truncation of singular values, while SVD-LLM Wang et al. (2025a) and its variant SVD-LLMv2 Wang et al. (2025b) proposed truncation-aware and optimized singular value selection strategies. In parallel, Basis Sharing Wang et al. (2024) explored cross-layer sharing of a common basis to further enhance compression efficiency by exploiting inter-layer redundancy. Despite their effectiveness, these methods are inherently constrained by the use of a single shared low-dimensional subspace, which may limit representational flexibility and prevent full exploitation of redundancies in LLM parameters.
In this work, we argue that moving beyond the low-rank weight decomposition is a promising new direction. Specifically, we propose to adopt dictionary learning, a well-established paradigm in signal and image processing Aharon et al. (2006); Elad (2010), and apply it to LLM compression. Unlike low-rank approximations that confine all columns of a weight matrix to a shared linear subspace, dictionary learning enables a richer union-of-subspaces representation: each column is approximated using only a sparse subset of atoms from a learned dictionary, allowing different columns to reside in distinct subspaces spanned by different atom combinations. Such a representation better accommodates heterogeneous features and introduces additional flexibility to reduce the overall approximation error.
To this end, we introduce CoSpaDi (Compression via Sparse Dictionary Learning), a training-free framework that applies dictionary learning to jointly estimate dictionaries and sparse coefficients used for compressing LLM weight matrices. Our contributions are as follows:
  • We introduce dictionary learning with sparse coding as a new paradigm for LLM compression, addressing the limitation of using a single invariant basis for the weight approximation which is inherent in SVD-based methods.
  • We demonstrate that CoSpaDi is effectively integrated with data-aware optimization, yielding state-of-the-art compression performance for a wide range of compression ratios, while being compatible with post-training quantization for further memory and latency gains.
  • Through extensive experiments, we show that our approach is effective in both per-layer and cross-layer (shared dictionary) compression scenarios, consistently outperforming regular SVD, ASVD Yuan et al. (2023), SVD-LLM Wang et al. (2025a), and Basis Sharing Wang et al. (2024) strategies as well as state-of-the-art structure-pruning methods Ma et al. (2023); Shopkhoev et al. (2025).
@@ 2 Related Work
Research on Transformer compression covers pruning, quantization, distillation, and matrix factorization. Below we provide a brief overview of the recent progress in these directions.
Early work on pruning showed substantial redundancy in deep nets Han et al. (2015); Frankle & Carbin (2019); Sanh et al. (2020). For LLMs, SparseGPT performs one-shot post-training pruning and achieves high sparsity with small quality drop Frantar & Alistarh (2023). Wanda reduces overhead further with simple activation-based rules Sun et al. (2023). LLM-Pruner removes blocks using gradient-guided criteria and recovers accuracy with brief adapter tuning Ma et al. (2023), while ReplaceMe substitutes multiple transformer blocks with a single linear transformation Shopkhoev et al. (2025).
Quantization reduces precision to reduce memory consumption and accelerate inference. For weight-only post-training quantization, LLM.int8 enables INT8 inference via vector-wise quantization with a path for outliers Dettmers et al. (2022); GPTQ/OPTQ minimize output error using Hessian to reach 3–4 bit weights Frantar et al. (2022; 2023); AWQ derives scales from activations and protects a small set of salient weights Lin et al. (2024), and SpQR stores outliers at higher precision to maintain quality at 3–4 bits Dettmers et al. (2024). For quantizing both weights and activations, SmoothQuant rebalances channel ranges to achieve accurate W8A8 across matrix multiplications Xiao et al. (2023), while QuaRot applies orthogonal rotations to suppress outliers, enabling end-to-end 4-bit inference including the KV cache Ashkboos et al. (2024).
The general framework on knowledge distillation by Hinton et al. and sequence-level distillation by Kim and Rush established strong baselines Hinton et al. (2015); Kim & Rush (2016). DistilBERT demonstrates task-agnostic compression for Transformers Sanh et al. (2019); TinyBERT uses a two-stage distillation procedure Jiao et al. (2020); MobileBERT introduces a teacher-guided compact architecture Sun et al. (2020); Patient-KD leverages multi-layer hints Sun et al. (2019); MiniLM and MiniLMv2 distill self-attention relations Wang et al. (2020; 2021). BERT-of-Theseus compresses models via progressive module replacement Xu et al. (2020). Orca and Orca 2 show that richer teacher signals such as explanations can improve reasoning in smaller students Mukherjee et al. (2023); Mitra et al. (2023).
Classic low-rank approximations reduce parameters and FLOPs with limited loss Denton et al. (2014). DRONE considers an objective related to the output activation for the weight approximation, Chen et al. (2021). FWSVD introduces Fisher-weighted reconstruction to emphasize important directions Hsu et al. (2022), while ASVD adapts truncation using activation-aware transforms and layer sensitivity Yuan et al. (2023). SVD-LLM utilizes truncation-aware whitening , while SVD-LLM v2 optimizes budget allocation across layers Wang et al. (2025a; b). Basis Sharing reuses one part of low-rank factorization across layers and learns layer-specific coefficients to exploit inter-layer redundancy Wang et al. (2024).
Dictionary learning and sparse coding. Dictionary learning factorizes original weight into dictionary and sparse codes, yielding a union-of-subspaces model with strong results in vision and image compression Engan et al. (1999); Aharon et al. (2006); Mairal et al. (2010); Gregor & LeCun (2010); Elad (2010). In NLP, GroupReduce explores block-wise low-rank/dictionary-style compression Chen et al. (2018), and tensorized Transformers factorize attention and embeddings Ma et al. (2019). Recent work also targets the KV-cache by learning universal dictionaries and decoding with OMP during inference Kim et al. (2024). Complementary to these directions, a recent work proposes cross-layer weight sharing via a matrix-based dictionary learning formulation for Transformer attention, directly exploiting inter-layer redundancy Zhussip et al. (2025). Cross-layer/shared-basis schemes in attention are thus closely related Wang et al. (2024). A comprehensive, training-free approach to weight-space dictionary learning for LLMs — at both per-layer and cross-layer levels — remains underexplored and is the focus of this work.
Figure 1: Left side: weight factorization methods using low-rank decomposition. Low-rank approximation decomposes a matrix into two dense matrices of lower rank. Right side: proposed CoSpaDi. A dictionary of $k$ atoms and a column-sparse coefficient matrix are employed. No restrictions on size of $k$ (undercomplete : $k<d_{1}$, complete: $k=d_{1}$ or overcomplete : $k>d_{1}$ dictionaries are possible), while sparsity is defined by $s$ non-zero elements per column of the coefficient matrix.
3 Proposed Method
In this section, we first provide a conceptual link between low-rank weight approximation and dictionary learning. Then, we introduce the proposed Sparse Dictionary Learning (SDL) strategy for LLM compression covering all practical aspects including data-aware SDL, dictionary-sharing among layers, compression ratios and inference complexity.
3.1 Low-Rank Weight Approximation
A widely adopted strategy to reduce the parameter count in LLMs is to approximate weight matrices $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ — representing the parameters of network layers — with matrices of reduced rank $r<\min(d_{1},d_{2})$. Such a low-rank approximation can be derived as follows:
$\tilde{\mathbf{W}}^{\star}=\operatorname*{arg\,min}_{\operatorname*{rank}\left(\tilde{\mathbf{W}}\right)=r}\left\|\mathbf{W}-\tilde{\mathbf{W}}\right\|_{F}.$ (1)
@@ 3.1 Low-Rank Weight Approximation
According to the Eckart–Young–Mirsky theorem Eckart & Young (1936), the orthogonal projection to the space of $r$-rank matrices admits an analytical solution. Specifically, if $\mathbf{W}$ admits the singular value decomposition $\mathbf{W}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\mathsf{T}}$, with $\mathbf{U}\in\mathbb{R}^{d_{1}\times k}$, $\bm{\Sigma}\in\mathbb{R}^{k\times k}$ and $\mathbf{V}\in\mathbb{R}^{d_{2}\times k}$ with $k=\min\left(d_{1},d_{2}\right)$, then the minimizer of Eq. (1) can be expressed as: $\tilde{\mathbf{W}}^{\star}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$, where $\mathbf{U}_{r}\in\mathbb{R}^{d_{1}\times r}$ contains the first $r$ left singular vectors, $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ holds the top-$r$ singular values, and $\mathbf{V}_{r}\in\mathbb{R}^{d_{2}\times r}$ contains the first $r$ right singular vectors of $\mathbf{W}$, respectively.
@@
There is a deep connection between this problem and the principal component analysis (PCA) Bishop & Nasrabadi (2006). Specifically, consider the weight matrix $\mathbf{W}$ as a collection of $d_{1}$-dimensional vectors $\mathbf{W}=\begin{bmatrix}\mathbf{w}_{1},&\ldots,\mathbf{w}_{d_{2}}\end{bmatrix}$, where $\bm{w}_{j}\in\mathbb{R}^{d_{1}}$, with $j=1,\ldots,d_{2}$. We seek to approximate each vector $\mathbf{w}_{j}$ as a linear combination of basis vectors spanning a lower-dimensional subspace of $\mathbb{R}^{d_{1}}$. The optimal basis and coefficients can be found by minimizing the total approximation error:
The optimal pair $\left(\mathbf{B}^{*},\mathbf{C}^{*}\right)$ is obtained by solving:
$\mathbf{B}^{*},\mathbf{C}^{*}=\operatorname*{arg\,min}_{\mathbf{B},\mathbf{C}}\mathcal{J}\,\,\mbox{s.t}\,\,\mathbf{B}^{\mathsf{T}}\mathbf{B}=\mathbf{I}.$ (3)
The solution is given by $\mathbf{B}^{*}=\mathbf{U}_{r}$ and $\mathbf{C}^{*}=\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$ (we refer to Appendix A.1 for the proof). This result highlights that in the low-rank factorization $\mathbf{W}\approx\mathbf{B}\mathbf{C}$, the matrix $\mathbf{B}$ corresponds to the basis of the shared $r$-dimensional subspace in which the columns of $\mathbf{W}$ approximately lie, and $\mathbf{C}$ contains their coordinate vectors (coefficients) with respect to that basis.
3.2 Sparse Dictionary Learning
Motivated by this interpretation, we propose an alternative strategy for compressing the weights $\mathbf{W}$, based on the sparse dictionary learning methodology. Specifically, rather than modeling each column of the weight matrix as a linear combination of basis vectors $\bm{b}_{i}$, we aim to learn a dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ and approximate each column $\mathbf{w}_{j}$ as a linear combination of dictionary atoms $\mathbf{d}_{i}\in\mathbb{R}^{d_{1}}\text{ with }i=1,\ldots,k$.
Although the two strategies appear similar, they differ in two key respects. First, in dictionary learning, unlike low-rank approximation, only a subset of atoms is used to represent each weight vector $\mathbf{w}_{j}$. This principle, known in the literature as sparse coding Lee et al. (2006), is motivated by the observation that learned dictionaries typically yield sparse representations: each signal (column) activates only a few atoms. In other words, instead of approximating the weight matrix $\mathbf{W}$ as the product of two dense matrices, the orthogonal basis $\mathbf{B}$ and the coefficient matrix $\mathbf{C}$, we approximate it as the product of a dense dictionary matrix $\mathbf{D}$ and a column-wise sparse coefficient matrix $\mathbf{S}$, where each column $\mathbf{s}_{j}$ indicates which atoms contribute to the representation of $\mathbf{w}_{j}$, and with what weights. Second, and of significant practical consequence, we do not impose orthogonality constraints on the dictionary atoms. This grants greater representational flexibility, allowing the model to adapt more effectively to the intrinsic structure of the weight matrix.
Finally, under the proposed strategy, rather than enforcing a shared low-dimensional subspace for all columns, each weight vector $\mathbf{w}_{j}$ is represented within its own low-dimensional subspace spanned by the atoms it activates. Consequently, the entire matrix $\mathbf{W}$ is modeled as a union of low-dimensional subspaces, introducing additional representational capacity and enabling lower approximation error compared to a single shared subspace.
Formally, the dictionary learning problem can be expressed as:
$\mathbf{D}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D},\mathbf{S}}\left\|\mathbf{W}-\mathbf{D}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2},$ (4)

### v3

# Denis Makhov1    Dmitriy Shopkhoev1,2 Magauiya Zhussip1    Ammar Ali1,2    Stamatios Lefkimmiatis1 Fundamental Research Center

@@ Abstract
Post-training compression of large language models (LLMs) often relies on low-rank weight approximations that represent each column of the weight matrix in a shared low-dimensional subspace. This strategy is computationally efficient but the underlying constraint can be overly rigid for heterogeneous projection weights and may incur avoidable accuracy loss. We propose CoSpaDi (Compression via Sparse Dictionary Learning), a training-free framework that replaces low-rank factorization with a structured sparse decomposition in which each weight matrix is represented as a dense dictionary multiplied by a column-sparse coefficient matrix. This yields a union-of-subspaces model: the columns of the weight matrix are represented as linear combinations of different subsets of dictionary atoms, improving expressiveness at a fixed parameter budget. CoSpaDi is calibration-guided: using a small calibration set, we optimize the factorization to minimize functional reconstruction error of layer outputs rather than weight-space error. An activation-derived Gram orthonormalization reformulates this data-aware objective into a standard dictionary learning problem on transformed weights, and we support both per-layer compression and cross-layer dictionary sharing within groups of similar projections. Across Llama and Qwen model families, CoSpaDi consistently improves the accuracy–compression and perplexity–compression trade-offs over state-of-the-art SVD-based baselines and strong structured pruning baselines at 20-40% compression ratios. The resulting structured sparsity enables sparse–dense computation and integrates with post-training quantization of the sparse coefficients.
@@ 1 Introduction
Large language models (LLMs) achieve strong performance across diverse tasks, from dialogue and instruction following (Brown et al., 2020; Achiam et al., 2023) to general-purpose reasoning (Touvron et al., 2023; Anil et al., 2023). Their effectiveness stems in part from their Transformer-based architectures that model long-range dependencies with attention (Vaswani et al., 2017; Devlin et al., 2019). At the same time, the scale that enables these capabilities makes LLMs expensive to store and run, creating a practical barrier to deployment on memory- and compute-constrained hardware.
A broad literature addresses post-training compression and acceleration of LLMs, spanning pruning, quantization, knowledge distillation, and structured weight parameterizations such as matrix factorization (Frankle & Carbin, 2019; Dettmers et al., 2022; Hinton et al., 2015; Denton et al., 2014). Among training-free approaches, matrix factorization is particularly attractive because it yields explicit low-parameter surrogates for large projection matrices, and in practice the predominant choice is truncated SVD and its activation-aware variants that use a small calibration set to guide which components to keep and how to scale them (Chen et al., 2021). Cross-layer extensions further reduce overhead by sharing a common low-dimensional subspace across groups of layers (Wang et al., 2025a). Despite strong results, these methods ultimately approximate each matrix within a single shared low-dimensional subspace; for heterogeneous Transformer projections, this constraint can be unnecessarily restrictive and motivates richer factorizations.
We study an alternative factorization family that relaxes this constraint. Instead of approximating a matrix with one global low-dimensional basis, we model it using a union of subspaces via sparse dictionary learning (Aharon et al., 2006; Elad, 2010): a learned dictionary of atoms is combined with column-sparse coefficients, allowing different columns to be reconstructed from different subsets of atoms. This representation is more flexible than a single shared subspace at the same parameter budget, and it is well aligned with the intuition that different output channels may depend on different latent features.
Building on this idea, we propose CoSpaDi (Compression via Sparse Dictionary Learning), a training-free compression framework for Transformer projections. CoSpaDi learns dictionaries and sparse codes to approximate pretrained weight matrices, and is data-aware: from a small calibration set we construct an activation-derived Gram orthonormalization that reformulates functional output reconstruction into a standard dictionary learning problem on a transformed (activation-weighted) weight matrices. The resulting factorization yields structured sparsity that can be paired with post-training quantization of the sparse coefficients, and it naturally supports both per-layer compression and cross-layer dictionary sharing for groups of related projections.
Contributions.
  • Beyond low-rank for post-training compression. We introduce sparse dictionary learning as a compression paradigm for LLM weight matrices, replacing the single-subspace constraint of SVD-based factorization with a union-of-subspaces representation.
  • Calibration-guided, training-free optimization. We integrate sparse dictionary learning with a data-aware objective via activation-derived Gram orthonormalization, yielding a tractable transformed problem that can be optimized efficiently with alternating sparse coding and dictionary updates, without gradient-based fine-tuning.
  • Empirical gains across settings. Across multiple Llama and Qwen models and a range of compression ratios, CoSpaDi improves the quality–compression trade-off over strong activation-aware SVD baselines (including SVD-LLM (Wang et al., 2025c) in the per-layer setting and Basis Sharing (Wang et al., 2025a) in the grouped setting), and is competitive with recent structured pruning methods (Ma et al., 2023; Shopkhoev et al., 2025).
@@ 2 Related Work
Post-training compression for Transformers is commonly organized around what is preserved (weights, activations, or end-to-end behavior) and what structure is imposed (sparsity, low precision, low rank, or shared components). We review the most relevant directions through this lens, and position CoSpaDi as a functional, calibration-guided method that imposes a union-of-subspaces structure via sparse dictionary learning.
Post-training pruning and practical sparsity.
A central theme in recent LLM compression is to optimize a compressed representation to match layer outputs on a small calibration set, rather than minimizing weight-space error. This perspective appears prominently in one-shot unstructured pruning methods such as SparseGPT (Frantar & Alistarh, 2023) and its simplified activation-driven variants (e.g., Wanda) (Sun et al., 2024), which induce high sparsity with limited degradation by explicitly targeting output preservation. However, unstructured sparsity does not automatically translate into practical compression or speedups: sparse storage formats typically require additional metadata (e.g., indices in CSR/CSC-like representations), which can substantially reduce effective memory gains at moderate sparsity, and irregular access patterns often make sparse matmuls memory-bound and kernel-limited (Han et al., 2015; Wang, 2020). In practice, inference acceleration usually benefits most from structured sparsity patterns and dedicated kernel/hardware support (e.g., 2:4 sparse tensor cores) (Mishra et al., 2021; NVIDIA, 2020). Complementary structured approaches remove or replace higher-level components (blocks/layers) using importance criteria and brief recovery steps (Ma et al., 2023; Shopkhoev et al., 2025). Our work follows the same calibration-guided spirit, but instead of removing weights, we seek a structured factorization that can represent heterogeneous columns more flexibly than a shared low-dimensional subspace.
Quantization and equivalent transformations.
Quantization reduces memory footprint and can improve throughput, but accuracy at low bit-width often hinges on handling outliers and reshaping activation/weight distributions. Weight-only PTQ methods include vector-wise quantization with outlier routing (Dettmers et al., 2022), Hessian-aware schemes such as GPTQ/OPTQ (Frantar et al., 2023a; b), and activation-aware scaling/outlier protection (Lin et al., 2024; Dettmers et al., 2024). For weight-and-activation quantization, methods such as SmoothQuant (Xiao et al., 2023) and rotation-based transforms (e.g., QuaRot) (Ashkboos et al., 2024) improve quantizability by rebalancing channel ranges or suppressing outliers. These techniques are complementary to structural compression: in our setting, the sparse coefficients produced by CoSpaDi can be quantized post hoc to further reduce memory and potentially improve practical efficiency.
Low-rank factorization and shared subspaces.
A widely used post-training alternative to pruning is matrix factorization, where each projection is approximated with a low-rank surrogate (Denton et al., 2014). Activation-aware and functional variants include DRONE (Chen et al., 2021), Fisher-weighted reconstruction (FWSVD) (Hsu et al., 2022), and adaptive truncation with activation transforms (ASVD) (Yuan et al., 2023). More recent SVD-LLM methods incorporate truncation-aware whitening and improved allocation strategies across layers (Wang et al., 2025c; b; Qinsi et al., 2025). In parallel, Basis Sharing (Wang et al., 2025a) ties factors across layers to exploit inter-layer redundancy. While effective, these approaches still represent each matrix in a single shared subspace (possibly partially shared across layers). Our method targets the same post-training regime, but replaces the single-subspace constraint with a union-of-subspaces representation enabled by sparse coding.
Dictionary learning and sparse representations.
Dictionary learning and sparse coding have a long history in signal processing and vision (Engan et al., 1999; Aharon et al., 2006; Mairal et al., 2009; Gregor & LeCun, 2010; Elad, 2010), where they provide expressive, parsimonious representations that adaptively select a small subset of atoms per example. In model compression, related ideas appear in structured decompositions such as GroupReduce (Chen et al., 2018) and tensorized factorization schemes (Ma et al., 2019). Recent work has also applied learned dictionaries to other Transformer components, e.g., KV-cache compression via sparse decoding (Kim et al., 2025), and cross-layer sharing formulations specialized to attention (Zhussip et al., 2025), which connects to broader shared-basis approaches (Wang et al., 2025a). In contrast, our focus is a training-free, calibration-guided dictionary learning framework for weight compression that (i) explicitly optimizes a functional objective through activation-aware Gram orthonormalization, and (ii) supports both per-layer compression and cross-layer dictionary sharing within a unified procedure.
Figure 1: Left: low-rank factorization represents $\mathbf{W}$ in a single $r$-dimensional subspace via two dense factors. Right: CoSpaDi represents $\mathbf{W}$ as $\mathbf{D}\mathbf{S}$ where $\mathbf{D}$ is a dictionary of $k$ atoms and $\mathbf{S}$ is column-sparse (at most $s$ nonzeros per column), yielding a union-of-subspaces model. Dictionaries may be undercomplete ($k<d_{1}$), complete ($k=d_{1}$), or overcomplete ($k>d_{1}$).
3 Method
We first formulate post-training compression objectives in weight space and activation space. We then reinterpret truncated SVD through the lens of PCA as a “basis times coefficients” model, and introduce sparse dictionary learning as a more expressive union-of-subspaces alternative. Finally, we describe CoSpaDi and its grouped extension, together with compression and complexity estimates.
3.1 Problem formulation: weight-space vs. activation-space objectives
Consider a pretrained linear projection with weight matrix $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ and input $\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$ (a batch of $N$ calibration activations). The layer output is $\mathbf{Z}=\mathbf{X}\mathbf{W}\in\mathbb{R}^{N\times d_{2}}$. Post-training compression replaces $\mathbf{W}$ by a structured approximation $\tilde{\mathbf{W}}$ that reduces storage and ideally improves inference efficiency.
A common starting point is weight-space reconstruction:
$\tilde{\mathbf{W}}^{\star}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}\in\mathcal{C}}\|\mathbf{W}-\tilde{\mathbf{W}}\|_{F}^{2},$ (1)
where $\mathcal{C}$ encodes a chosen structure (e.g., rank-$r$, sparsity, quantization, or a parametric factorization).
For LLM projections, however, preserving weights is often a weak proxy for preserving model behavior. A more directly relevant goal is activation-space (functional) reconstruction on calibration data:
$\tilde{\mathbf{W}}^{\star}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}\in\mathcal{C}}\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\|_{F}^{2}.$ (2)
Objective (2) is used (implicitly or explicitly) across many calibration-guided post-training methods, and motivates the activation-aware design of CoSpaDi.
3.2 Truncated SVD as PCA: basis $\times$ coefficients
Low-rank approximation chooses $\mathcal{C}=\{\tilde{\mathbf{W}}:\mathrm{rank}(\tilde{\mathbf{W}})\leq r\}$ and solves (1):
$\tilde{\mathbf{W}}^{\mathrm{LR}}=\operatorname*{arg\,min}_{\mathrm{rank}(\tilde{\mathbf{W}})\leq r}\|\mathbf{W}-\tilde{\mathbf{W}}\|_{F}^{2}.$ (3)
@@ 3.2 Truncated SVD as PCA: basis $\times$ coefficients
According to the Eckart–Young–Mirsky theorem (Eckart & Young, 1936), the orthogonal projection to the space of $r$-rank matrices admits an analytical solution. Specifically, if $\mathbf{W}$ admits the singular value decomposition $\mathbf{W}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\mathsf{T}}$, with $\mathbf{U}\in\mathbb{R}^{d_{1}\times k}$, $\bm{\Sigma}\in\mathbb{R}^{k\times k}$ and $\mathbf{V}\in\mathbb{R}^{d_{2}\times k}$ with $k=\min\left(d_{1},d_{2}\right)$, then the minimizer of Eq. (3) can be expressed as: $\tilde{\mathbf{W}}^{\mathrm{LR}}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}$, where $\mathbf{U}_{r}\in\mathbb{R}^{d_{1}\times r}$ contains the first $r$ left singular vectors, $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ holds the top-$r$ singular values, and $\mathbf{V}_{r}\in\mathbb{R}^{d_{2}\times r}$ contains the first $r$ right singular vectors of $\mathbf{W}$, respectively.
@@
There is a deep connection between this problem and the principal component analysis (PCA) (Bishop & Nasrabadi, 2006). Specifically, consider the weight matrix $\mathbf{W}$ as a collection of $d_{1}$-dimensional vectors $\mathbf{W}=\begin{bmatrix}\mathbf{w}_{1},&\ldots,\mathbf{w}_{d_{2}}\end{bmatrix}$, where $\bm{w}_{j}\in\mathbb{R}^{d_{1}}$, with $j=1,\ldots,d_{2}$. We seek to approximate each vector $\mathbf{w}_{j}$ as a linear combination of basis vectors spanning a lower-dimensional subspace of $\mathbb{R}^{d_{1}}$. The optimal basis $\mathbf{B}\in\mathbb{R}^{d_{1}\times r}$ and coefficients $\mathbf{C}\in\mathbb{R}^{r\times d_{2}}$ can be found by minimizing the total approximation error:
$\min_{\mathbf{B},\mathbf{C}}\ \|\mathbf{W}-\mathbf{B}\mathbf{C}\|_{F}^{2}\quad\text{s.t.}\ \mathbf{B}^{\top}\mathbf{B}=\mathbf{I}.$ (4)
The optimum is $\mathbf{B}^{\star}=\mathbf{U}_{r}$ and $\mathbf{C}^{\star}=\bm{\Sigma}_{r}\mathbf{V}_{r}^{\top}$ (see Appendix A.1 for the proof). Therefore, truncated SVD realizes a single shared $r$-dimensional subspace: every column $\mathbf{w}_{j}$ is expressed in the same basis $\mathbf{B}$ via dense coordinates (the $j$-th column of $\mathbf{C}$).
Although Eq. (3) minimizes reconstruction error in weight space, several works (e.g., (Chen et al., 2021)) note that LLM projection matrices often lack strong intrinsic low-rank structure, so directly approximating $\mathbf{W}$ can lead to noticeable degradation. A common alternative is to optimize functional reconstruction on calibration activations, motivated by the hypothesis that layer inputs lie in a lower-dimensional subspace and that preserving the induced outputs is more relevant than preserving weights in isolation. Concretely, this leads to minimizing $\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\|_{F}^{2}$ for calibration inputs $\mathbf{X}$ (Chen et al., 2021; Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2025c). In this setting, the optimal low-rank approximation can still be obtained in closed form after applying an activation-derived Gram orthonormalization; we provide the full derivation and solution in Appendix A.2.
3.3 Sparse dictionary learning: union of subspaces
Dictionary learning keeps the “basis times coefficients” form but changes the modeling assumptions. We replace the orthonormal basis $\mathbf{B}$ by a (generally non-orthogonal) dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ and enforce column sparsity in the coefficient matrix $\mathbf{S}\in\mathbb{R}^{k\times d_{2}}$:
$\min_{\mathbf{D},\mathbf{S}}\ \|\mathbf{W}-\mathbf{D}\mathbf{S}\|_{F}^{2}\quad\text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s,\ \forall j,$ (5)